OpenStudy (anonymous):

Can someone solve this problem: A wireless phone company charges a monthly fee of $8.95 and a cost per minute of use. One customer was charged $20.35 for 95 minutes. How much will another customer pay for 243 minutes? Is the answer $38.11?

OpenStudy (anonymous):

how much is per min

OpenStudy (anonymous):

yes the answer is 38.11

OpenStudy (anonymous):

it's pretty simple really, just subtract the monthly fee from the 20.35

OpenStudy (anonymous):

then divide that by 95 mins to get $0.12 per min

OpenStudy (anonymous):

$20.35=95x + $8.95 - 8.95 - 8.95 11.40 = 95x (divide both sides by 95) .12 = x THEN Y= 243(.12) + 8.95 Y= 29.16 + 8.95 Y= 38.11

OpenStudy (anonymous):

^^^^ Is that correct?

OpenStudy (anonymous):

so the cust who used 243 mins spent $29.16 plus the monthly fee

OpenStudy (anonymous):

yea, I think that way works too. it was just faster for me to work it in my head the way I did it.

OpenStudy (anonymous):

Yeah. Thanks for the clarification.. Can I ask you something else?

OpenStudy (anonymous):

you can ask, not sure if I'll have an answer :-)

OpenStudy (anonymous):

How would you explain my work in words..?

OpenStudy (anonymous):

algebraaaaaaaa

OpenStudy (anonymous):

uhmm.... well you said the billed amount is equal to the number of mins plus the monthly fee

OpenStudy (anonymous):

subtracted the monthly fee from both sides to eliminate it from the equation

OpenStudy (anonymous):

and divided the amount charged/min by the number of mins to get the cost/min of 12 cents

OpenStudy (anonymous):

and then you said, total mins times 12 cents plus the monthly fee of 8.95 and voila

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

no problem!